OpenStudy (anonymous):

\[\int\limits_{}^{}x^{3}dx/\sqrt{2 + x^{4}}\] How do I solve this? I assume I use trigonometric substitution but when I do I'm still left with a very ugly integral.

OpenStudy (turingtest):

how about the sub\[u=2+x ^4\]??

OpenStudy (anonymous):

u = 2 + x^(4) du/3 = x^(3)dx 1/(u)^(1/2) makes sense thanks